OpenStudy (anonymous):

solve the equation for x. k+1=z(x+p)

OpenStudy (anonymous):

\[k + 1 = zx + zp\]\[k + 1 - zp = zx\]\[\frac{k+1 - zp}{z}\]

OpenStudy (anonymous):

solve for x not k...

OpenStudy (anonymous):

??

OpenStudy (anonymous):

is the answer k+1-p/z

OpenStudy (anonymous):

I got K+1over z-p=x I don't know if it's right